# The Milky Way in context: The formation of galactic discs and chemical sequences from a cosmological perspective

### v1

@@ 4.3 Fuelling from the CGM
This is a result that is backed up by idealised simulations in Khoperskov et al. (2021), in which the low-$\alpha$ sequence forms from a mixture of metal-poor CGM and enriched gases that were ejected by stellar feedback during the formation of the high-$\alpha$ sequence. Furthermore, they find that the formation of two chemical sequences is not dependent on any one particular merger history or on radial migration.
Other fully cosmological simulations also highlight the importance of the CGM. Analysis of the FIRE-2 cosmological simulations in Parul et al. (2025) showed that bi-modal sequences arise when metal-poor gas accretes onto the outskirts of the disc, owing to both merger donation and the CGM. However, their realisation with a high fraction of merger-donated gases (36 per cent) is only weakly bi-modal, and they conclude that merger-donated gases do not necessarily lead to the development of a distinct chemical sequence.
Investigations into the higher resolution versions of auriga in Grand et al. (2018) show that late-time minor mergers can penetrate the hot gas halo and deliver cold metal-poor gas directly onto the galaxy. While our own analysis shows that the merger gas neither dominates the low-$\alpha$ sequence nor reduces the metallicity by more than $\sim 0.3\,$dex, it nonetheless plays an important role in triggering the growth of the gas disc. The merger may also incite the condensation of surrounding metal-poor CGM gas onto the disc, leading to an indirect dilution effect.
In the vintergatan cosmological simulations, a disc with bi-modal chemical sequences forms in the presence of a GSE-like merger (Agertz et al., 2021; Renaud et al., 2021). In particular, their low-$\alpha$ sequence exhibits a ‘reset’ in its metallicity. A detailed investigation revealed that the gas donated by the GSE-like merger was unimportant for both the formation of the low-$\alpha$ sequence and the metallicity dilution. Instead, they owe to an accreting metal-poor and cold gas filament that is initially misaligned with respect to the gas disc. The GSE-like merger is nonetheless essential because its interaction with this accreting gas filament acts to excite star formation within it. Similarly, we find that mergers are important for triggering star formation, even if their merger mass ratios are less than $1:100$. This is a result also echoed in Parul et al. (2025).
@@ 4.4 Our interpretation for the MW history
The MW began to form early, with an inside-out and top-down formation of the stellar disc. These early stars formed during a period of intense SFR from a turbulent ISM. Whilst the bulk of the high-$\alpha$ sequence formed at this time, there may have been sub-dominant populations of low-$\alpha$ stars forming in dense gas clumps.
@@
At around 10 Gyr ago, the GSE merger accretes onto the proto-MW. Following reasoning in the literature (e.g. Kruijssen et al., 2019, 2020; Lane et al., 2023), this was a minor merger with a mass ratio $<1:4$. This event initiates a combination of physical processes (inward funnelling of CGM gas, donation of fresh cold gas, gas shocks) which conspire to induce a central star formation burst. This burst bolsters the number of stars in the high-$\alpha$ sequence, but also acts to deplete the star-forming gas, and is followed by a reduction of the SFR that lasts $\mathcal{O}1\,$Gyr. The GSE has other implications on the fledgling MW disc, kinematically heating its stars and shaping what is now the thick disc (e.g. see Grand et al. 2020).
@@
Assuming a variety of masses and metallicities for the GSE progenitor galaxy and contemporary proto-MW, we show in Appendix A that the GSE merger would have been incapable of driving a substantial metallicity ‘reset’ between the metal-rich end of the high-$\alpha$ sequence and the metal-poor end of the low-$\alpha$ sequence. Instead, continuous feeding from the CGM is responsible for maintaining a metal-poor zone at the height and radial extremities of the gas disc (see Figure 11). The metal-poor ends of the low-$\alpha$ sequence is comprised from stars that form out of these gases.
@@
With a relatively quiescent history thereafter, the MW disc becomes thinner and more stable, and the SFR gradually declines as the gas density in the centre dwindles. This is shown in the decline of the local SFR from around 6 Gyrs ago in Ruiz-Lara et al. (2020), and around 2.5 Gyrs ago in Fernández-Alvar et al. (2025), though with significant fluctuations in both cases. The outcome is the development of a radial gradient in the [$\alpha$/Fe] ratio, which grows from the inside-out as the enrichment from SNe Ia starts to dominate over the enrichment from SNe II. As we show in Section 3.3.4, this gradient can be imprinted on the low-$\alpha$ sequence as a downward ‘slant’ (a negative relation between [Fe/H] and [Mg/Fe], and see also Mason et al. 2024).
@@
Several key elements are still missing from this interpretation. First, there is a lack of high-quality observational analogues of the MW at redshifts $z>2$. Current surveys continue to face challenges in resolving galaxies at the necessary mass scales at such early epochs. However, gravitational lensing techniques (Jones et al., 2023) offer a promising avenue for overcoming these limitations by magnifying faint and distant systems, though further progress is needed to fully exploit their potential. It is also crucial to identify observational evidence of dense, star-forming gas clumps and the specific environments in which they emerge. Finally, improved age statistics for stellar populations in the MW is essential, though upcoming missions like PLATO (Rauer et al., 2025) may provide some relief on this front.
@@ 5 Conclusions
  • Half of the thirty auriga simulations exhibit bi-modal distribution in their [Mg/Fe] distribution, with a further two having multi-modal distributions. Regardless of the presence of a bi-modality, high-$\alpha$ and low-$\alpha$ populations usually adhere to the same typical bulk properties – those being hotter kinematics (with higher $\sigma_{z}$ by an average of $33\,\rm{km}\,\rm{s}^{-1}$), slower rotational velocities (by an average of $32\,\rm{km}\,\rm{s}^{-1}$), and older ages for the high-$\alpha$ stars (by an average of 4.6 Gyrs).
@@
  • Singular merger events, in particular those that are GSE-like, do not dilute the metallicity of the CGM by more than $\approx 0.3\,$dex, and usually by even less. The gas donated by GSE-like mergers forms only a minority of subsequent disc stars. The only exceptions are in the cases of very massive mergers (merger mass ratios of $>1:2$, and total masses of $>2\times 10^{10}$), though these are ruled out for the MW. Therefore, the GSE gas was not responsible for the overlapping metallicities of the high and low-$\alpha$ sequences, and we have reinforced this with empirical estimates. Other merger events such as kraken will have been just as unimportant (see also arguments in Orkney et al., 2022).
@@
  • The break between two chemical sequences is driven by a temporary quenching or reduction of the SFR. This can be due to multiple reasons, including gas exhaustion, post-merger quenching, or a violent disruption of the gas disc. As such, massive mergers such as the GSE can potentially contribute to the development of chemical bi-modalities due to their merger-induced starbursts and the following quiescence. Other mechanisms may also be important, such as AGN-induced shutdowns or bar formation (e.g. Beane 2025; Beane et al. 2024), though these may themselves be triggered by a merger event. Disentangling these processes is beyond the scope of this work.
@@
Our results highlight the importance of obtaining high-quality observational data for MW analogues at higher redshifts ($z\geq 2$). The James Webb Space Telescope (JWST) has recently begun probing such galaxies (e.g. Mowla et al., 2024; Tan et al., 2024b, a), and surveys like JADES may already be detecting MW analogues as early as $z\approx 8.2$ (Rusta et al., 2024).
As observational data for M31 and other Local Volume galaxies will continue to improve, understanding the formation of chemical abundance patterns beyond the MW will become increasingly important. Our investigation revealed a broad diversity in the chemical evolution of MW-mass disc galaxies, which are highly sensitive to factors such as overall mass assembly, star formation history, and even the specifics of individual merger events. To better constrain the physical and dynamical processes that set chemical trends in MW-like galaxies, detailed extragalactic observations within the Local Volume, as well as the CGM of high-redshift galaxies, will prove essential in the era of Extremely Large Telescopes.
@@ 6 Acknowledgements
CL & MO acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 852839). RJJG is supported by an STFC Ernest Rutherford Fellowship (ST/W003643/1).
@@
The auriga simulation data is publicly available to download via the Globus platform as described in section 4 of Grand et al. (2017). Any novel post-processed data is available upon request.
Appendix A Estimating the gas dilution in the MW due to the GSE
Model $M_{\star}$ [M] $M_{\rm gas}$ [M] $M_{\rm DM}$ [M] Gas [Fe/H]
@@ Appendix A Estimating the gas dilution in the MW due to the GSE
Table 1: The model properties for the proto-MW and GSE merger at a time of 10 Gyr ago or $z\approx 2$. Sources are as follows: a) Limberg et al. (2022), b) Lane et al. (2023), c) Helmi et al. (2018). The methods used to derive these model parameters are described in the main text of Section A.
@@
We estimate the gas metallicity by considering that the stars born at the metal-rich end of the MW high-$\alpha$ sequence will reflect the ambient gas metallicity before the GSE infall. From visual inspection of the figures in Queiroz et al. (2020), we take $\rm{[Fe/H]}=0.00\pm 0.05$ as a rough guideline. Then, the amount of dilution required to reach the metal-poor end of the low-$\alpha$ sequence would be around -0.5 dex. This is far less than what is presented in the chemical evolution models of (Spitoni et al., 2019, 2021, 2023), but we consider it as a conservative minimum required dilution.
@@
An alternative estimate of the metallicity can be obtained using the stellar mass–metallicity relation from Kirby et al. (2013) (given by their equation 4, hereafter Kirby13 MZR), which gives $\rm{[Fe/H]}(\rm{proto\text{-}MW})=-0.58^{+0.27}_{-0.16}$. Although this relation is calibrated on galaxy observations at $z=0$, it nonetheless captures the spread and median of observed stellar metallicities within the high-$\alpha$ sequence. However, it is not suitable for our purposes, as we are interested in the chemical evolution stage near the end of the high-$\alpha$ sequence where metallicities approach $\rm{[Fe/H]}\approx 0$.
@@
For the properties of the GSE progenitor galaxy, we construct three models based upon literature values. We assume the gas metallicity matches the stellar metallicity. Average stellar enrichment lags behind the gas enrichment, and so this choice should be considered as a low estimate of the gas metallicity. We calculate the gas mass assuming high gas fractions of 0.7-0.9, reflecting the high gas fractions expected for ancient dwarf galaxies (i.e. Popping et al., 2012). Once again, the halo masses are approximated from the Moster13 SMHM. Our models are described as follows:
@@
  • GSE 1: We take the GSE stellar metallicity value of $\rm{[Fe/H]}=-1.22\pm 0.23$ from Limberg et al. (2022) and find the corresponding stellar mass from the Kirby13 MZR. We note that Limberg et al. (2022) also quote their own stellar mass estimate of $\sim 1.3\times 10^{9}\,\rm{M}_{\odot}$, but we determine that this is likely too high because the entire MW stellar halo is only $\sim 1.4\times 10^{9}\,\rm{M}_{\odot}$ (Deason et al., 2019).
@@
  • GSE 2: We take the GSE stellar mass value of $M_{\star}=1.45^{+0.92}_{-0.51}\times 10^{8}\,\rm{M}_{\odot}$ from Lane et al. (2023) and find the corresponding stellar metallicity from the inverted Kirby13 MZR.
@@
  • GSE 3: Here, we take the direct GSE stellar mass and metallicity values reported in Helmi et al. (2018) as $M_{\star}=6\times 10^{8}\,\rm{M}_{\odot}$ and $\rm{[Fe/H]}=-1.6$, and note that these values are significantly offset from the Kirby13 MZR. This model can be considered as a particularly high-mass and metal-poor estimate for the GSE, thereby maximising any resulting dilution.
@@
The uncertainties on the parameters in Table 1 are either taken from their respective sources, or are approximated as the ±$\pm\sigma$ percentiles from a Monte Carlo method using $10^{6}$ samples. Symmetric uncertainties are sampled from a normal distribution, whereas the upper and lower bounds for asymmetric uncertainties are calculated independently with a two-piece normal distribution provided by the twopiece python package. The distributions in $f_{\rm gas}$ are sampled from a uniform distribution. The Kirby13 MZR and Moster13 SMHM have a further 0.17 and 0.15 dex scatter respectively, which we include in addition to the parametric uncertainties of each relation. The GSE parameters from Helmi et al. (2018) do not include uncertainties, and so we caution the reader that the true uncertainties will be greater than those shown here.
@@
We calculate the gas dilution by averaging the metallicities of the GSE model and the proto-MW, weighting by their respective gas masses. It is conceivable that the GSE gas did not mix homogeneously with the proto-MW gas, and so we choose to calculate the dilution over a range of proto-MW gas masses in Figure 15. There, the vertical black line indicates our estimate for the true gas mass of the proto-MW from Table 1, with associated uncertainty represented as a shaded band. This figure shows that the most likely outcome is for a dilution magnitude of $<0.2$, very similar to the maximum dilution magnitudes that we estimated for auriga in panels (b) and (c) of Figure 7, but far from the required 0.5 dex needed in the MW. The dilution magnitude remains relatively low even in cases where the GSE gas has a greater mass than the proto-MW gas, and this is due to the logarithmic nature of the [Fe/H] ratio.
@@
The GSE may have deposited its gas around the edge of the proto-MW gas disc, where the gas densities are lower and higher dilutions can be achieved more easily. An example of this scenario is described in auriga by Grand et al. (2018), for a late-time minor merger in the higher-resolution version of Au-23. This merger has a wide pericentre passage of $>40\,$kpc, which enables its gas to be gradually stripped away and accreted onto the edge of the gas disc. This is a less likely possibility for the GSE, which must have had a highly radial infall trajectory with a low pericentre.
Our results are consistent with the study by Palla et al. (2020). They model the gas contribution from the GSE, calibrated to match the abundances reported in Helmi et al. (2018). The results indicate that the merger supplied approximately 15 per cent of the total gas during the second-infall epoch, remarkably close to the fractions shown in our Figure 8 in the case of a $1:8$ mass ratio merger. They also find a minimal metal dilution, no matter the assumed accretion time or the chemical composition of the merger gas.

### v2

@@ 4.3 Fuelling from the CGM
This is a result that is backed up by idealised simulations in khoperskov2021, in which the low-$\alpha$ sequence forms from a mixture of metal-poor CGM and enriched gases that were ejected by stellar feedback during the formation of the high-$\alpha$ sequence. Furthermore, they find that the formation of two chemical sequences is not dependent on any one particular merger history or on radial migration.
Other fully cosmological simulations also highlight the importance of the CGM. Analysis of the FIRE-2 cosmological simulations in parul2025 showed that bi-modal sequences arise when metal-poor gas accretes onto the outskirts of the disc, owing to both merger donation and the CGM. However, their realisation with a high fraction of merger-donated gases (36 per cent) is only weakly bi-modal, and they conclude that merger-donated gases do not necessarily lead to the development of a distinct chemical sequence.
Investigations into the higher resolution versions of auriga in grand2018 show that late-time minor mergers can penetrate the hot gas halo and deliver cold metal-poor gas directly onto the galaxy. While our own analysis shows that the merger gas neither dominates the low-$\alpha$ sequence nor reduces the metallicity by more than $\sim 0.3\,$dex, it nonetheless plays an important role in triggering the growth of the gas disc. The merger may also incite the condensation of surrounding metal-poor CGM gas onto the disc, leading to an indirect dilution effect.
In the vintergatan cosmological simulations, a disc with bi-modal chemical sequences forms in the presence of a GSE-like merger (agertz2021; agertz2021b). In particular, their low-$\alpha$ sequence exhibits a ‘reset’ in its metallicity. A detailed investigation revealed that the gas donated by the GSE-like merger was unimportant for both the formation of the low-$\alpha$ sequence and the metallicity dilution. Instead, they owe to an accreting metal-poor and cold gas filament that is initially misaligned with respect to the gas disc. The GSE-like merger is nonetheless essential because its interaction with this accreting gas filament acts to excite star formation within it. Similarly, we find that mergers are important for triggering star formation, even if their merger mass ratios are less than $1:100$. This is a result also echoed in parul2025.
@@ 4.4 Our interpretation for the MW history
The MW began to form early, with an inside-out and top-down formation of the stellar disc. These early stars formed during a period of intense SFR from a turbulent ISM. Whilst the bulk of the high-$\alpha$ sequence formed at this time, there may have been sub-dominant populations of low-$\alpha$ stars forming in dense gas clumps. The proto-galaxy would have undergone a series of high mass-ratio mergers at this time, perhaps including the speculative kraken (i.e. Kruijssen2019; massari2019). These mergers contribute gaseous fuel that would have been important for developing the early gas disc and initiating star formation in the high-$\alpha$ sequence, but would have been relatively unimportant in terms of their direct chemical impact (see orkney2022 and arguments in Section 3.3).
@@
At around 10 Gyr ago, the GSE merger accretes onto the proto-MW. Following reasoning in the literature (e.g. Kruijssen2019; kruijssen2020; lane2023), this was a minor merger with a mass ratio <1:4<1:4. This event initiates a combination of physical processes (inward funnelling of CGM gas, donation of fresh cold gas, gas shocks) which conspire to induce a central star formation burst. This burst bolsters the number of stars in the high-$\alpha$ sequence, but also acts to deplete the star-forming gas, and is followed by a reduction of the SFR that lasts $\mathcal{O}1\,$Gyr. The GSE has other implications on the fledgling MW disc, kinematically heating its stars and shaping what is now the thick disc (e.g. see grand2020).
@@
Assuming a variety of masses and metallicities for the GSE progenitor galaxy and contemporary proto-MW, we show in Appendix B that the GSE merger would have been incapable of driving a substantial metallicity ‘reset’ between the metal-rich end of the high-$\alpha$ sequence and the metal-poor end of the low-$\alpha$ sequence. Instead, continuous feeding from the CGM is responsible for maintaining a metal-poor zone at the height and radial extremities of the gas disc (see Figure 11). The metal-poor ends of the low-$\alpha$ sequence is comprised from stars that form out of these gases.
@@
With a relatively quiescent history thereafter, the MW disc becomes thinner and more stable, and the SFR gradually declines as the gas density in the centre dwindles. This is shown in the decline of the local SFR from around 6 Gyrs ago in ruizlara2020, and around 2.5 Gyrs ago in Fernandez2025, though with significant fluctuations in both cases. The outcome is the development of a radial gradient in the $\alpha$/Fe] ratio, which grows from the inside-out as the enrichment from SNe Ia starts to dominate over the enrichment from SNe II. As we show in Section 3.3.4, this gradient can be imprinted on the low-$\alpha$ sequence as a downward ‘slant’ (a negative relation between [Fe/H] and [Mg/Fe], and see also mason2024).
@@
Several key elements are still missing from this interpretation. First, there is a lack of high-quality observational analogues of the MW at redshifts $z>2$. Current surveys continue to face challenges in resolving galaxies at the necessary mass scales at such early epochs. However, gravitational lensing techniques (jones2023) offer a promising avenue for overcoming these limitations by magnifying faint and distant systems, though further progress is needed to fully exploit their potential. It is also crucial to identify observational evidence of dense, star-forming gas clumps and the specific environments in which they emerge. Finally, improved age statistics for stellar populations in the MW is essential, though upcoming missions like PLATO (PLATO) may provide some relief on this front.
Our investigation is also relevant to M31, for which early spectroscopic results suggest a single chemical sequence similar to the MW high-$\alpha$ disc (escala2020; nidever2024). Inspection of Figure 4 in nidever2024 reveals a narrow chemical track that spans from high-$\alpha$, low-metallicity to low-$\alpha$, high-metallicity stars. Compared to our analysis, this distribution closely resembles realisations that form an early thin gaseous disc (e.g., Au-2 and Au-24). This appears to conflict with observational evidence from dust geometry indicating that the dominant disc component of M31 has a thickness comparable to the MW thick disc (dalcanton2023). However, these kinematics may be shaped by a possible major merger 2-4,Gyr ago (see arguments in hammer2018; Bhattacharya2024; tsakonas2025).
@@ 5 Conclusions
  • Half of the thirty auriga simulations exhibit bi-modal distribution in their [Mg/Fe] distribution, with a further two having multi-modal distributions. Regardless of the presence of a bi-modality, high-$\alpha$ and low-$\alpha$ populations usually adhere to the same typical bulk properties — those being hotter kinematics (with higher $\sigma_{z}$ by an average of $33\,\rm{km}\,\rm{s}^{-1}$), slower rotational velocities (by an average of $32\,\rm{km}\,\rm{s}^{-1}$), and older ages for the high-$\alpha$ stars (by an average of 4.6 Gyrs).
@@
  • Singular merger events, in particular those that are GSE-like, do not dilute the metallicity of the CGM by more than $\approx 0.3\,$dex, and usually by even less. The gas donated by GSE-like mergers forms only a minority of subsequent disc stars. The only exceptions are in the cases of very massive mergers (merger mass ratios of >1:2>1:2, and total masses of $2\times 10^{10}$), though these are ruled out for the MW. Therefore, the GSE gas was not responsible for the overlapping metallicities of the high and low-$\alpha$ sequences, and we have reinforced this with empirical estimates. Other merger events such as kraken will have been just as unimportant, even if they were of higher mass ratio (see also arguments in orkney2022).
@@
  • The break between two chemical sequences is driven by a temporary quenching or reduction of the SFR. This can be due to multiple reasons, including gas exhaustion, post-merger quenching, or a violent disruption of the gas disc. As such, massive mergers such as the GSE can potentially contribute to the development of chemical bi-modalities due to their merger-induced starbursts and the following quiescence. Other mechanisms may also be important, such as AGN-induced shutdowns or bar formation (e.g. beane2024a; beane2024b), though these may themselves be triggered by a merger event. Disentangling these processes is beyond the scope of this work.
@@
Our results highlight the importance of obtaining high-quality observational data for MW analogues at higher redshifts ($z\geq 2$). The James Webb Space Telescope (JWST) has recently begun probing such galaxies (e.g. mowla2024; tan2024; tan2024b), and surveys like JADES (JWST Advanced Deep Extragalactic Survey) may already be detecting MW analogues as early as $z\approx 8.2$ (rusta2024).
Early spectroscopic studies of M31 suggest that its chemistry may not be bi-modal (escala2020; nidever2024). As observations of other Local Volume galaxies continue to improve, understanding the formation of chemical abundance patterns beyond the MW will become increasingly important. Our investigation revealed a broad diversity in the chemical evolution of MW-mass disc galaxies, which are highly sensitive to factors such as overall mass assembly, star formation history, and even the specifics of individual merger events. To better constrain the physical and dynamical processes that set chemical trends in MW-like galaxies, detailed extragalactic observations within the Local Volume, as well as the CGM of high-redshift galaxies, will prove essential in the era of Extremely Large Telescopes (see hammer2021; padovani2023).
@@ 6 Acknowledgements
We thank the anonymous referee for their thoughtful comments, which have greatly contributed to this manuscript. This work has used resources from the MareNostrum 4 supercomputer at the Barcelona Supercomputing Center (BSC). Much of our analysis was performed on the Virgo supercomputer at the Max Planck Computing and Data Facility (MPCDF). Further analysis and simulation work was performed on the NYX supercomputer at the Universität de Barcelona (ICCUB). CL & MO acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 852839). CL also acknowledges funding from the Agence Nationale de la Recherche (ANR project ANR-24-CPJ1-0160-01). The author acknowledges financial support from the CEX2024-001451-M grant funded by MICIU/AEI/10.13039/501100011033. RJJG is supported by an STFC Ernest Rutherford Fellowship (ST/W003643/1).
@@
The auriga simulation data is publicly available to download via the Globus platform as described in section 4 of Auriga. Any novel post-processed data is available upon request.
Appendix A Full star formation history plot
Figure 15: The same as in Figure 4, but now for the entire auriga suite.
We include star formation histories for the full auriga suite in Figure 15.
Appendix B Estimating the gas dilution in the MW due to the GSE
Model $_{\star} [M] M$$_{\rm gas} [M$] $M_{\rm DM}$ [M] Gas [Fe/H]
@@ Appendix B Estimating the gas dilution in the MW due to the GSE
Table 1: The model properties for the proto-MW and GSE merger at a time of 10 Gyr ago or $z\approx 2$. Sources are as follows: a) limberg2022, b) lane2023, c) helmi2018. The methods used to derive these model parameters are described in the main text of Section B.
@@
We estimate the gas metallicity by considering that the stars born at the metal-rich end of the MW high-$\alpha$ sequence will reflect the ambient gas metallicity before the GSE infall. From visual inspection of the figures in queiroz2020, we take $\rm{[Fe/H]}=0.00\pm 0.05$ as a rough guideline. Then, the amount of dilution required to reach the metal-poor end of the low-$\alpha$ sequence would be around -0.5 dex. This is far less than what is presented in the chemical evolution models of (spitoni2019; spitoni2021; spitoni2023), but we consider it as a conservative minimum required dilution.
@@
An alternative estimate of the metallicity can be obtained using the stellar mass–metallicity relation from kirby2013 (given by their equation 4, hereafter Kirby13 MZR), which gives $\rm{[Fe/H]}(\rm{proto\text{-}MW})=-0.58^{+0.27}_{-0.16}$. Although this relation is calibrated on galaxy observations at $z=0$, it nonetheless captures the spread and median of observed stellar metallicities within the high-$\alpha$ sequence. However, it is not suitable for our purposes, as we are interested in the chemical evolution stage near the end of the high-$\alpha$ sequence where metallicities approach $\rm{[Fe/H]}\approx 0$.
@@
For the properties of the GSE progenitor galaxy, we construct three models based upon literature values. We assume the gas metallicity matches the stellar metallicity. Average stellar enrichment lags behind the gas enrichment, and so this choice should be considered as a low estimate of the gas metallicity. We calculate the gas mass assuming high gas fractions of 0.7-0.9, reflecting the high gas fractions expected for ancient dwarf galaxies (i.e. popping2012). Once again, the halo masses are approximated from the Moster13 SMHM. Our models are described as follows:
@@
  • GSE 1: We take the GSE stellar metallicity value of $\rm{[Fe/H]}=-1.22\pm 0.23$ from limberg2022 and find the corresponding stellar mass from the Kirby13 MZR. We note that limberg2022 also quote their own stellar mass estimate of $\sim 1.3\times 10^{9}\,\rm{M}_{\odot}$, but we determine that this is likely too high because the entire MW stellar halo is only $\sim 1.4\times 10^{9}\,\rm{M}_{\odot}$ (deason2019).
@@
  • GSE 2: We take the GSE stellar mass value of $M_{\star}=1.45^{+0.92}_{-0.51}\times 10^{8}\,\rm{M}_{\odot}$ from lane2023 and find the corresponding stellar metallicity from the inverted Kirby13 MZR.
@@
  • GSE 3: Here, we take the direct GSE stellar mass and metallicity values reported in helmi2018 as $M_{\star}=6\times 10^{8}\,\rm{M}_{\odot}$ and $\rm{[Fe/H]}=-1.6$, and note that these values are significantly offset from the Kirby13 MZR. This model can be considered as a particularly high-mass and metal-poor estimate for the GSE, thereby maximising any resulting dilution.
@@
The uncertainties on the parameters in Table 1 are either taken from their respective sources, or are approximated as the $\pm\sigma$ percentiles from a Monte Carlo method using $10^{6}$ samples. Symmetric uncertainties are sampled from a normal distribution, whereas the upper and lower bounds for asymmetric uncertainties are calculated independently with a two-piece normal distribution provided by the twopiece python package. The distributions in $f_{\rm gas}$ are sampled from a uniform distribution. The Kirby13 MZR and Moster13 SMHM have a further 0.17 and 0.15 dex scatter respectively, which we include in addition to the parametric uncertainties of each relation. The GSE parameters from helmi2018 do not include uncertainties, and so we caution the reader that the true uncertainties will be greater than those shown here.
@@
We calculate the gas dilution by averaging the metallicities of the GSE model and the proto-MW, weighting by their respective gas masses. It is conceivable that the GSE gas did not mix homogeneously with the proto-MW gas, and so we choose to calculate the dilution over a range of proto-MW gas masses in Figure 16. There, the vertical black line indicates our estimate for the true gas mass of the proto-MW from Table 1, with associated uncertainty represented as a shaded band. This figure shows that the most likely outcome is for a dilution magnitude of <0.2<0.2, very similar to the maximum dilution magnitudes that we estimated for auriga in panels (b) and (c) of Figure 7, but far from the required 0.5 dex needed in the MW. The dilution magnitude remains relatively low even in cases where the GSE gas has a greater mass than the proto-MW gas, and this is due to the logarithmic nature of the [Fe/H] ratio.
@@
The GSE may have deposited its gas around the edge of the proto-MW gas disc, where the gas densities are lower and higher dilutions can be achieved more easily. An example of this scenario is described in auriga by grand2018, for a late-time minor merger in the higher-resolution version of Au-23. This merger has a wide pericentre passage of >40>40\,kpc, which enables its gas to be gradually stripped away and accreted onto the edge of the gas disc. This is a less likely possibility for the GSE, which must have had a highly radial infall trajectory with a low pericentre.
Our results are consistent with the study by palla2020. They model the gas contribution from the GSE, calibrated to match the abundances reported in helmi2018. The results indicate that the merger supplied approximately 15 per cent of the total gas during the second-infall epoch, remarkably close to the fractions shown in our Figure 8 in the case of a $1:8$ mass ratio merger. They also find a minimal metal dilution, no matter the assumed accretion time or the chemical composition of the merger gas.